hhanan:

What is the equation of the sequence given? 2500, -500, 100, -20...

jhonyy9:

how you get from 2500 the - 500 ?

smackz:

tbh i forgot this . but ask @Laylalyssa @boredfr @yogirlkaylie @nevaehhh

smackz:

or johnny9 the papas

yogirlkaylie:

huh

jhonyy9:

or from -500 the 100 ?

jhonyy9:

or from 100 the -20

CripQUEZZ:

Ok so look as it u got -20 to 100 So how would u get to that? Divide 100 by -20 what will u get then u just multiply by the number u get

hhanan:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 or from -500 the 100 ? \(\color{#0cbb34}{\text{End of Quote}}\) add 600

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hhanan \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 or from -500 the 100 ? \(\color{#0cbb34}{\text{End of Quote}}\) add 600 \(\color{#0cbb34}{\text{End of Quote}}\) No It’s a division problem bc ur going backwards so jus divide -500 by 100

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hhanan \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 or from -500 the 100 ? \(\color{#0cbb34}{\text{End of Quote}}\) add 600 \(\color{#0cbb34}{\text{End of Quote}}\) by divide it

mxddi3:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hhanan \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 or from -500 the 100 ? \(\color{#0cbb34}{\text{End of Quote}}\) add 600 \(\color{#0cbb34}{\text{End of Quote}}\) yes, in that one scenario, yes, but it doesn't work for all the other scenarios. try another way (:

hhanan:

-5

CripQUEZZ:

Correct

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hhanan -5 \(\color{#0cbb34}{\text{End of Quote}}\) Yes good job

mxddi3:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hhanan -5 \(\color{#0cbb34}{\text{End of Quote}}\) yes, we are dividing by -5. 2500/-5=-500 -500/-5=100 and so on.

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mxddi3 \(\color{#0cbb34}{\text{Originally Posted by}}\) @hhanan -5 \(\color{#0cbb34}{\text{End of Quote}}\) yes, we are dividing by -5. 2500/-5=-500 -500/-5=100 and so on. \(\color{#0cbb34}{\text{End of Quote}}\) Yupp

hhanan:

so i put -5 as the equation of the sequence?

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hhanan so i put -5 as the equation of the sequence? \(\color{#0cbb34}{\text{End of Quote}}\) Yes the equation is x/-5 = y so like 2500/-5 = -500

hhanan:

ty

CripQUEZZ:

Np